OpenStudy (anonymous):

The manager of bulk foods establisher sells trail mix for $9 per pound and premium cashews for $15 per pound. The manager wishes to make a 108-pound trail mix-cashews mixture that will sell for $14 per pound. How any pounds of cashews should be used?

OpenStudy (shane_b):

Just make the system of equations using the information that was given: Let x = number of pounds of trail mix and y = number of pounds of cashews: 9x + 15y = 108*14 x + y = 108 Now solve for x and y using substitution...

OpenStudy (anonymous):

54.

OpenStudy (shane_b):

You should have ended up with x=18 lbs trailmix and y = 90 lbs cashews.